# VAGUE: Visual Contexts Clarify Ambiguous Expressions

### v2

@@ B.5 Model Instruction for Direct and Indirect prompts
Fig. H8 is the prompt used to generate direct expressions. Also, Fig. H9 is a prompt that generates the correct answer for multiple choice by understanding the intention based on this direct expression(Correct). Fig. H10 is the prompt that generates indirect expressions based on direct expressions and their intended meaning. Additionally, it produces superficially misinterpreted intentions (SU).
@@ B.6 Human Rating for Direct and Indirect prompts
This section details the human rating and filtering process. Human annotators are instructed to follow the scoring guidelines provided in the tables in Figs. H13 and H14. To facilitate high-quality annotations efficiently, we use Label Studio (https://labelstud.io/). Figure H13 illustrates the UI for rating direct expressions, while Fig. H14 shows the UI for selecting and rating indirect expressions.
@@ B.7 Model Instruction for Counterfactual Choices
Fig. H11 is the prompt used to generate incorrect answer choices by creating fake captions that aligns with the indirect expressions but is inconsistent with the direct expressions and combining fake captions with the indirect expressions to derive the plausible intention (FS). Fig. H12 is the prompt that generates incorrect answer choices by introducing objects not present in the image, intentionally misrepresenting the intended meaning of the indirect expression (NE).
@@ Appendix C Baseline Models
To ensure broad coverage of existing multimodal language models (MLLMs), we evaluate eight open-source models with varying parameter sizes alongside two closed-source models. For the open-source models, we use Phi3.5-Vision-Instruct (4B) [27], optimized for concise instruction-following tasks, providing robust text-image alignment in low-parameter settings. LLaVA Onevision (7B) [22] focuses on high-resolution image interpretation, enhancing multimodal dialogue through refined attention mechanisms. Qwen2.5-VL-Instruct (7B) [40] uses advanced vision-language pretraining to handle diverse image-based queries and textual instructions. InternVL-2.5-MPO (8B, 26B) [6] is designed for multi-purpose optimizations, supporting enhanced multimodal reasoning. Idefics2 (8B) [21] adopts a compact architecture for efficient training, emphasizing domain-specific image understanding and textual generation. LLaVA NeXT Vicuna (13B) [23] employs an improved vision encoder and refined instruction tuning for enhanced commonsense reasoning. Ovis2 (16B) [26] excels in image captioning and inference, driven by robust textual grounding and visual alignment. Among proprietary models, GPT-4o [29] demonstrates advanced language comprehension paired with visual perception for nuanced multimodal interactions. Gemini 1.5 Pro [11] integrates high-resolution vision processing with a powerful language model, delivering refined instruction-following and cross-domain reasoning.
@@ D.3 Qualitative Example
Fig. H4 shows the results of free-form answering in the VLM, SM, and LM settings using the InternVL-2.5-8B-MPO model. It shows that the underlying intention behind the indirect expression in the given image is well preserved, and the generated responses across different settings are highly similar.
@@ Appendix E Details of Human Evaluation
Appendix F Model Instruction for Experiments
@@
Fig. H15, Fig. H16, Fig. H17 show the prompts used for multiple-choice question experiments under VLM, SM, and LM settings.
@@
Fig. H18, Fig. H19, Fig. H20 show the prompts used for free-form answering experiments under VLM, SM, and LM settings.
@@
Additionally, Fig. H24, Fig. H22 are prompts that we use for zeroshot chain-of-thought experiments with Socratics Models, while Fig. H23, Fig. H21 are those with Visual-Language Models.
Appendix G Full Results
@@
Table H3 and Table H4 present the complete experimental results conducted in this paper for the VAGUE-VCR and VAGUE-Ego4D datasets, respectively. The total number of items in VAGUE-VCR is 1,144, and in VAGUE-Ego4D, it is 533. However, the item count in the valid count column does not always match these totals. This discrepancy occurs when the model responds with ’I don’t know’ or refuses to answer. When calculating accuracy, we treat such cases as incorrect and divide the number of correct responses by the total number of items in each dataset.
Appendix H Full structure of VAGUE
Fig. H25 shows the structure of our benchmark dataset, VAGUE.
@@ Appendix H Full structure of VAGUE
Figure H3: This table presents two evaluation criteria for direct expressions and two for indirect expressions, along with descriptions and corresponding bad and good cases. The examples in the bad and good cases are derived from human ratings based on the given criteria.
@@
Figure H4: This figure shows the results generated by InternVL-2.5-8B-MPO for indirect expressions and free-form responses across VLM, SM, and LM settings.
@@
Figure H5: These are examples of an image along with its corresponding generated direct expression and multiple-choice question set.
@@
Figure H6: These are examples of an image along with its corresponding generated direct expression and multiple-choice question set.
@@
Figure H7: These are examples of an image along with its corresponding generated direct expression and multiple-choice question set.
@@
Figure H8: This prompt selects one of the list of entities and generates a direct request to a person in the image. It also generates a triple solution and generates a caption for the scene.
@@
Figure H9: This prompt takes an action, object, and direct expression as input and outputs the true intention behind the direct expression.
@@
Figure H10: This prompt takes a direct expression, true intention, and caption as input to generate an indirect expression that conveys the intended meaning. Additionally, it produces a superficially interpreted version of the indirect expression.
@@
Figure H11: This prompt uses a fake caption and an indirect expression to derive the most plausible intention that does not align with the true intention of direct expression.
@@
Figure H12: This prompt takes an indirect expression and an interpretation of the indirect expression as input to generate an incorrect intention based on a nonexistent object in the image.
@@
Figure H13: The actual interface used for human rating of direct expressions. It includes the two evaluation criteria for direct expressions and detailed scoring guidelines on a 1 to 5 scale.
@@
Figure H14: The actual interface used for human rating of indirect expressions. It includes the two evaluation criteria for indirect expressions and detailed scoring guidelines on a 1 to 5 scale.
@@
Figure H15: The following prompt is used for the VLM setting in the multiple-choice question task. The model receives an image, an utterance, and a set of answer choices as input and selects the most appropriate answer.
@@
Figure H16: The following prompt is used for the SM setting in the multiple-choice question task. The model first generates a caption for the image. Then, it receives the caption, an utterance, and a set of answer choices as input and selects the most appropriate answer.
@@
Figure H17: The following prompt is used for the LM setting in the multiple-choice question task. The model receives an utterance and a set of answer choices as input and selects the most appropriate answer.
@@
Figure H18: The following prompt is used for the VLM setting in the free-form answering task. The model receives an image, an utterance as input and it tasked with inferring the underlying intention.
@@
Figure H19: The following prompt is used for the SM setting in the free-form answering task. The model first generates a caption for the image. Then, it receives the caption and an utterance as input and it tasked with inferring the underlying intention.
@@
Figure H20: The following prompt is used for the LM setting in the free-form answering task. The model receives an utterance as input and it tasked with inferring the underlying intention.
@@
Figure H21: The following prompt is used for the VLM Chain of Thought setting in the multiple-choice question task. The model receives an image, an utterance, and a set of answer choices as input and selects the most appropriate answer by thinking step by step.
@@
Figure H22: The following prompt is used for the SM Chain of Thought setting in the multiple-choice question task. The model first generates a caption for the image. Then, it receives the caption, an utterance, and a set of answer choices as input and selects the most appropriate answer by thinking step by step.
@@
Figure H23: The following prompt is used for the VLM Chain of Thought setting in the free-form question task. The model receives an image, an utterance as inputs and it tasked with inferring the underlying intention by thinking step by step.
@@
Figure H24: The following prompt is used for the SM Chain of Thought setting in the free-form question task. The model first generates a caption for the image. Then, it receives the caption and an utterance as inputs and it tasked with inferring the underlying intention by thinking step by step.
@@
Accuracy(%) Incorrect Count Selection Count Valid Count Bert F1 BLEU(1gram) Valid Count
@@
Table H3: The overall results table for the VAGUE-VCR dataset. Experiments are conducted on both Multiple Choice Questions and Free-Form Answering, measuring results across three settings for each model: VLM, SM, and LM. For GPT-4o and Gemini 1.5 Pro, CoT reasoning is additionally applied in the VLM and SM settings.
@@
Accuracy(%) Incorrect Count Selection Count Valid Count Bert F1 BLEU(1gram) Valid Count
@@
Table H4: The overall results table for the VAGUE-Ego4D dataset. Experiments are conducted on both Multiple Choice Questions and Free-Form Answering, measuring results across three settings for each model: VLM, SM, and LM. For GPT-4o and Gemini 1.5 Pro, CoT reasoning is additionally applied in the VLM and SM settings.
Figure H25: We show the structure using a sample from our benchmark dataset, VAGUE. VAGUE consists of an image name, direct expression, indirect expression, triplet solution, multiple-choice set, meta data containing various information about the image, and a fake caption.

### v3

@@ B.5 Model Instruction for Direct and Indirect prompts
Fig. J8 is the prompt used to generate direct expressions. Also, Fig. J9 is a prompt that generates the correct answer for multiple choice by understanding the intention based on this direct expression(Correct). Fig. J10 is the prompt that generates indirect expressions based on direct expressions and their intended meaning. Additionally, it produces superficially misinterpreted intentions (SU).
@@ B.6 Human Rating for Direct and Indirect prompts
This section details the human rating and filtering process. Human annotators are instructed to follow the scoring guidelines provided in the tables in Figs. J13 and J14. To facilitate high-quality annotations efficiently, we use Label Studio (https://labelstud.io/). Figure J13 illustrates the UI for rating direct expressions, while Fig. J14 shows the UI for selecting and rating indirect expressions.
@@ B.7 Model Instruction for Counterfactual Choices
Fig. J11 is the prompt used to generate incorrect answer choices by creating fake captions that aligns with the indirect expressions but is inconsistent with the direct expressions and combining fake captions with the indirect expressions to derive the plausible intention (FS). Fig. J12 is the prompt that generates incorrect answer choices by introducing objects not present in the image, intentionally misrepresenting the intended meaning of the indirect expression (NE).
@@ Appendix C Baseline Models
To ensure broad coverage of existing multimodal language models (MLLMs), we evaluate ten open-source models with varying parameter sizes alongside two closed-source models. For the open-source models, we use Phi3.5-Vision-Instruct (4B) [29], optimized for concise instruction-following tasks, providing robust text-image alignment in low-parameter settings. LLaVA Onevision (7B) [23] focuses on high-resolution image interpretation, enhancing multimodal dialogue through refined attention mechanisms. Qwen2.5-VL-Instruct (7B, 72B) [43] uses advanced vision-language pretraining to handle diverse image-based queries and textual instructions. InternVL-2.5-MPO (8B, 26B) [6] is designed for multi-purpose optimizations, supporting enhanced multimodal reasoning. InternVL-3 (38B) [42] provides better performance in various tasks than in the previously released models. Idefics2 (8B) [22] adopts a compact architecture for efficient training, emphasizing domain-specific image understanding and textual generation. LLaVA NeXT Vicuna (13B) [24] employs an improved vision encoder and refined instruction tuning for enhanced commonsense reasoning. Ovis2 (16B) [28] excels in image captioning and inference, driven by robust textual grounding and visual alignment. Among proprietary models, GPT-4o [31] demonstrates advanced language comprehension paired with visual perception for nuanced multimodal interactions. Gemini 1.5 Pro [12] integrates high-resolution vision processing with a powerful language model, delivering refined instruction-following and cross-domain reasoning.
@@ D.3 Qualitative Example
Fig. J4 shows the results of free-form answering in the VLM, SM, and LM settings using the InternVL-2.5-8B-MPO model. It shows that the underlying intention behind the indirect expression in the given image is well preserved, and the generated responses across different settings are highly similar.
@@ Appendix E Details of Human Evaluation
Appendix F Why does SM setting outperform in proprietary models?
[…]
[] Models Self captioning GPT-4o captioning Phi3.5-Vision-Instruct (4b) 34.0 37.9 ($\uparrow$ 3.9) LLaVA-Onevision (7b) 29.4 36.5 ($\uparrow$ 7.1) LLaVA-NeXT-vicuna (13B) 36.3 41.7 ($\uparrow$ 5.4) InternVL-2.5-MPO (26B) 50.6 50.8 ($\uparrow$ 0.2)
Table F3: $\uparrow$ denotes MCQ performance boost in SM when switching to GPT-4o-generated captions.
The rationale benind the superior SM performance of proprietary models lies in their reduced vision-related (FS, NE) errors. We elaborate this in two points: 1) If captions lack necessary detail, VLMs naturally excel. Tab. F3 implies proprietary models have better captions, by showing a performance boost when leveraging GPT-4o’s captions in smaller open-source models within the SM setting. 2) If the caption somehow attains parity of detail (Practically, length constraints preclude full nuance capture.) with the image, text tokens handle visual information more effectively than image tokens [25]. Indeed, we observed in CoT reasoning that VLMs lean more on the speaker’s rationale while SMs take a more scene‐focused approach
Appendix G CoT Open-source
Model Type Acc (%) Incorrect count
FS SU NE
Phi3.5-Vision-Instruct (4B) SM 34.0 292 678 137
SM+CoT 31.5 ($\downarrow$ 2.5) 291 603 148
VLM 44.8 266 501 158
VLM+CoT 36.6 ($\downarrow$ 8.2) 295 510 154
LLaVA-Onevision (7B) SM 29.4 215 885 84
SM+CoT 28.0 ($\downarrow$ 1.4) 291 795 101
VLM 43.1 169 727 58
VLM+CoT 33.1 ($\downarrow$ 10.0) 187 736 76
LLaVA-NeXT-vicuna (13B) SM 36.3 228 730 111
SM+CoT 32.6 ($\downarrow$ 3.7) 259 598 149
VLM 48.4 206 564 96
VLM+CoT 38.3 ($\downarrow$ 10.1) 206 541 139
InternVL-2.5-MPO (26B) SM 50.6 209 530 89
SM+CoT 45.1 ($\downarrow$ 5.5) 213 584 106
VLM 65.3 147 377 58
VLM+CoT 58.1 ($\downarrow$ 7.2) 17 440 67
InternVL-3 (38B) SM 47.3 259 478 136
SM+CoT 54.3 ($\uparrow$ 7.0) 166 359 102
VLM 62.4 153 374 102
VLM+CoT 52.2 ($\downarrow$ 10.2) 211 387 107
Qwen2.5-VL-Instruct (72B) SM 56.8 175 457 92
SM+CoT 55.8 ($\downarrow$ 1.0) 213 428 100
VLM 72.8 142 236 78
VLM+CoT 69.9 ($\downarrow$ 2.9) 154 272 78
Table G4: Result of Chain-of-Thought (CoT) experiments on open-source models, in both SM and VLM settings. $\uparrow$ and $\downarrow$ indicate an increase and decrease in accuracy when zero-shot CoT is applied.
We provide CoT ablations on open-source models and compare them to our SM and VLM baselines. Tab. G4 shows that CoT slightly degrades SM in the most cases, while VLMs experience a much bigger drop in their performance, increasing all types of errors (FS, SU, NE) together. Consistent degradation shows open-source models lag proprietary ones in deep reasoning. We believe the sharper drop in VLMs stems from conditioning dilution [10] due to longer generation sequences.
Appendix H Model Instruction for Experiments
@@
Fig. J15, Fig. J16, Fig. J17 show the prompts used for multiple-choice question experiments under VLM, SM, and LM settings.
@@
Fig. J18, Fig. J19, Fig. J20 show the prompts used for free-form answering experiments under VLM, SM, and LM settings.
@@
Additionally, Fig. J24, Fig. J22 are prompts that we use for zeroshot chain-of-thought experiments with Socratics Models, while Fig. J23, Fig. J21 are those with Visual-Language Models.
Appendix I Full Results
@@
Table J5 and Table J6 present the complete experimental results conducted in this paper for the VAGUE-VCR and VAGUE-Ego4D datasets, respectively. The total number of items in VAGUE-VCR is 1,144, and in VAGUE-Ego4D, it is 533. However, the item count in the valid count column does not always match these totals. This discrepancy occurs when the model responds with ’I don’t know’ or refuses to answer. When calculating accuracy, we treat such cases as incorrect and divide the number of correct responses by the total number of items in each dataset.
Appendix J Full structure of VAGUE
Fig. J25 shows the structure of our benchmark dataset, VAGUE.
@@
Figure J3: This table presents two evaluation criteria for direct expressions and two for indirect expressions, along with descriptions and corresponding bad and good cases. The examples in the bad and good cases are derived from human ratings based on the given criteria.
@@
Figure J4: This figure shows the results generated by InternVL-2.5-8B-MPO for indirect expressions and free-form responses across VLM, SM, and LM settings.
@@
Figure J5: These are examples of an image along with its corresponding generated direct expression and multiple-choice question set.
@@
Figure J6: These are examples of an image along with its corresponding generated direct expression and multiple-choice question set.
@@
Figure J7: These are examples of an image along with its corresponding generated direct expression and multiple-choice question set.
@@
Figure J8: This prompt selects one of the list of entities and generates a direct request to a person in the image. It also generates a triple solution and generates a caption for the scene.
@@
Figure J9: This prompt takes an action, object, and direct expression as input and outputs the true intention behind the direct expression.
@@
Figure J10: This prompt takes a direct expression, true intention, and caption as input to generate an indirect expression that conveys the intended meaning. Additionally, it produces a superficially interpreted version of the indirect expression.
@@
Figure J11: This prompt uses a fake caption and an indirect expression to derive the most plausible intention that does not align with the true intention of direct expression.
@@
Figure J12: This prompt takes an indirect expression and an interpretation of the indirect expression as input to generate an incorrect intention based on a nonexistent object in the image.
@@
Figure J13: The actual interface used for human rating of direct expressions. It includes the two evaluation criteria for direct expressions and detailed scoring guidelines on a 1 to 5 scale.
@@
Figure J14: The actual interface used for human rating of indirect expressions. It includes the two evaluation criteria for indirect expressions and detailed scoring guidelines on a 1 to 5 scale.
@@
Figure J15: The following prompt is used for the VLM setting in the multiple-choice question task. The model receives an image, an utterance, and a set of answer choices as input and selects the most appropriate answer.
@@
Figure J16: The following prompt is used for the SM setting in the multiple-choice question task. The model first generates a caption for the image. Then, it receives the caption, an utterance, and a set of answer choices as input and selects the most appropriate answer.
@@
Figure J17: The following prompt is used for the LM setting in the multiple-choice question task. The model receives an utterance and a set of answer choices as input and selects the most appropriate answer.
@@
Figure J18: The following prompt is used for the VLM setting in the free-form answering task. The model receives an image, an utterance as input and it tasked with inferring the underlying intention.
@@
Figure J19: The following prompt is used for the SM setting in the free-form answering task. The model first generates a caption for the image. Then, it receives the caption and an utterance as input and it tasked with inferring the underlying intention.
@@
Figure J20: The following prompt is used for the LM setting in the free-form answering task. The model receives an utterance as input and it tasked with inferring the underlying intention.
@@
Figure J21: The following prompt is used for the VLM Chain of Thought setting in the multiple-choice question task. The model receives an image, an utterance, and a set of answer choices as input and selects the most appropriate answer by thinking step by step.
@@
Figure J22: The following prompt is used for the SM Chain of Thought setting in the multiple-choice question task. The model first generates a caption for the image. Then, it receives the caption, an utterance, and a set of answer choices as input and selects the most appropriate answer by thinking step by step.
@@
Figure J23: The following prompt is used for the VLM Chain of Thought setting in the free-form question task. The model receives an image, an utterance as inputs and it tasked with inferring the underlying intention by thinking step by step.
@@
Figure J24: The following prompt is used for the SM Chain of Thought setting in the free-form question task. The model first generates a caption for the image. Then, it receives the caption and an utterance as inputs and it tasked with inferring the underlying intention by thinking step by step.
@@
Accuracy(%) Incorrect Count Correct Valid Count Bert F1 BLEU(1gram) Valid Count
@@
Table J5: The overall results table for the VAGUE-VCR dataset. Experiments are conducted on both Multiple Choice Questions and Free-Form Answering, measuring results across three settings for each model: VLM, SM, and LM. For GPT-4o and Gemini 1.5 Pro, CoT reasoning is additionally applied in the VLM and SM settings.
@@
Accuracy(%) Incorrect Count Correct Valid Count Bert F1 BLEU(1gram) Valid Count
@@
Table J6: The overall results table for the VAGUE-Ego4D dataset. Experiments are conducted on both Multiple Choice Questions and Free-Form Answering, measuring results across three settings for each model: VLM, SM, and LM. For GPT-4o and Gemini 1.5 Pro, CoT reasoning is additionally applied in the VLM and SM settings.
Figure J25: We show the structure using a sample from our benchmark dataset, VAGUE. VAGUE consists of an image name, direct expression, indirect expression, triplet solution, multiple-choice set, meta data containing various information about the image, and a fake caption.